OpenStudy (anonymous):

let f(x)=1/x.find a number b so that the avg rate of change of f(x) on the interval [-b,b] is the value 1/2...how do i even set this up??

OpenStudy (cruffo):

\[ARC = \frac{f(x_2)-f(x_1)}{x_2-x_1}\] \[= \frac{f(b)-f(-b)}{b-(-b)}\] \[= \frac{\frac{1}{b}-\frac{1}{-b}}{b+b}\] \[= \frac{\frac{1}{b}+\frac{1}{b}}{2b}\] \[= \frac{\frac{2}{b}}{2b}\] \[=\frac{2}{b} \cdot \frac{1}{2b}\] \[=\frac{1}{b^2}\] We need the above fraction to be 1/2 so that means b^2 = 2, or b = sqrt(2).

OpenStudy (anonymous):

wouldn't it be 2b times 2 not 2 times 1on ur 2nd 2 last step? (dont u have to flip the fraction when u go from divide to mult.?

OpenStudy (anonymous):

i believe that is the step im confused on..

OpenStudy (anonymous):

ohhh, nvm got it.

OpenStudy (cruffo):

cool...